# Algorithms for Reconstructing B Cell Lineages in the Presence of Context-Dependent Somatic Hypermutation

### v1

@@ Abstract
We introduce a method for approximating posterior probabilities of phylogenetic trees and reconstructing ancestral sequences under models of sequence evolution with site-dependence, where standard phylogenetic likelihood computations (pruning) fail. Our approach uses a combined data-augmentation and importance sampling scheme. A key advantage of our approach is the ability to leverage existing highly optimized phylogenetic software. We apply our approach to the reconstruction of B cell receptor affinity maturation lineages from high-throughput repertoire sequencing data and evaluate the impact of incorporating site-dependence on the reconstruction accuracy of both trees and ancestral sequences. We show that accounting for context-dependence during inference always improves the estimates of both ancestral sequences and lineage trees on simulated datasets. We also examine the impact of incorporating priors based on VDJ recombination models, and find that they significantly improve ancestral sequence reconstruction in germline-encoded regions, but increase errors in non-templated nucleotides. We propose a modified, piecewise prior to address this demonstrate that it improves empirical reconstruction accuracy. We apply our approach to the analysis of the HIV broadly neutralizing antibodies DH270 and CH235 which are important targets of current vaccine design efforts.
@@ 1 Introduction
To prevent infectious diseases, the immune system must recognize and clear foreign pathogens capable of host infection. The adaptive immune system does so using B cells with receptors specifically tailored to bind pathogens. Facing a near-infinite variety of pathogen protein surfaces, B cells rely on two diversification methods: 1) initial gene segment shuffling (VDJ recombination) to enable weak binding to the diversity of antigens encountered, and 2) a refinement step (affinity maturation) that substantially improves binding and BCR-antigen recognition specificity, accomplished via somatic hypermutation (SHM). Understanding of this process and design of new vaccines is greatly aided by computational and statistical analysis of data from high-throughput B cell repertoire sequencing. A key challenge is accurate inference of BCR lineages and unmutated common ancestors, and significant progress has been made on these tools [kepler_reconstructing_2013, Matsen2016, Marcou2018, dhar_bayesian_2020]. However, existing methods rely on conventional “independent-site" mutation models from molecular phylogenetics and unlike mutational processes elsewhere in the genome, these models do not adequately capture the SHM process, which is mediated by the enzyme activation induced cytidine deaminase (AID). Critically, AID activity is sequence-context dependent, targeting sites in certain favored motifs. Unfortunately, incorporation of context-dependent mutation models makes standard phylogenetic calculations intractable. Here we develop algorithms for reconstructing B cell lineages under context-dependent SHM models. We build on recent work on approximating pairwise sequence evolution probabilities [Mathews:2023b, Mathews:2025a, Mathews:2025b], extending these in non-trivial ways to trees using Monte Carlo methods of data augmentation, importance sampling, and sequential Monte Carlo. Notably, our approach can be implemented directly using existing phylogenetics software.
@@ 2.1 BnAbs, Sequential Vaccine Design, and B Cell Lineages
The goal of vaccination is to induce antibodies – secreted versions of B cell receptors (BCRs) – that neutralize pathogens and protect against infection. In infection, successive rounds of mutation and selection for antigen affinity (affinity maturation) enables B cells to recognize pathogens with specificity. However, rapidly mutating viruses (HIV, influenza, SARS-CoV-2) escape this B cell response by viral diversification. It is thus desired to elicit broadly neutralizing antibodies (bnAbs), but these are rarely elicited due to a variety of factors [Walker_Burton_2018]: low precursor frequencies in the human B cell receptor repertoire [Haynes_35962033, Jardine_27013733], highly mutated, and containing essential mutations that are highly improbable, creating acquisition bottlenecks [wiehe_functional_2018, Saunders_Wiehe_2019, Mathews:2023b]. Traditional vaccine design strategies have proven ineffective at eliciting bnAbs, motivating new strategies using known bnAbs as templates to design immunogens to guide B cell maturation the bnAb [Haynes_2012, Jardine_2015, Haynes_35962033, Steichen_2019, Kwong_29768174]. Sequential prime-boosting aims to first induce low-frequency bnAb precursors with a priming immunogen, then mature them with additional boosting immunogens, and relies heavily on inferring the precursor BCR sequence (the unmutated common ancestor (UCA)) from observed, clonally-related sequences [kepler_reconstructing_2013, Haynes_2012]. More generally reconstruction of B cell lineage - the set of affinity maturation pathways for clone members - provides insight into maturation bottlenecks and also precursors to serve as intermediate vaccine targets. Existing lineage reconstruction methods such as Cloanalyst [kepler_reconstructing_2013], IgOR [Marcou2018], Partis [Matsen2016] and LinearHam [dhar_bayesian_2020] combine phylogenetic inference with models of VDJ recombination, but follow the vast majority of molecular phylogenetics in assuming that sequence positions evolve independently; this assumption is critical to tractability of phylogenetic inference. While appealing in their simplicity and computational convenience, these independent-site models fail to capture the critical context-dependence of SHM in affinity maturation.
@@ 2.2 Affinity Maturation of B Cell Receptors: Context-Dependent Somatic Hypermutation
S5F model and ARMADiLLO: We previously developed ARMADiLLO [wiehe_functional_2018] for simulating SHM using context-dependent mutation probabilities from the S5F model [Yaari2013], which captures context-dependence of SHM using context-dependent site-mutation and substitution probabilities derived from NGS data [Yaari2013] for the central nucleotide of 1,024 5-mer motifs, given its (2 upstream and 2 downstream) neighboring bases. The model parameterizes mutation probabilities as a product of mutability scores and substitution probabilities. ARMADiLLO has been used to forward-simulate SMH and as a null model to test for selection [wiehe_functional_2018, Tang:2023]. Unfortunately, this more realistic SHM model cannot be used in existing lineage reconstruction algorithms [kepler_reconstructing_2013, Marcou2018, Matsen2016, dhar_bayesian_2020], as context-dependent mutation models directly violate the independent-site assumption that makes standard phylogenetic calculations (pruning) tractable. Since we might expect improved SHM models to provide improved reconstruction of UCAs and lineages, it is of great interest to develop methods for B cell lineage reconstruction which can handle context-dependent mutation models such as S5F/ARMADiLLO.
@@ 3 Background
3.1 Standard Models of DNA Evolution   Let $\mathbf{x}=(x_{1},\ldots,x_{\ell})$ and $\mathbf{y}=(y_{1},\ldots,y_{\ell})$ denote two nucleotide sequences of length $\ell$, where $x_{i},y_{i}\in\mathscr{A}$ for $\mathscr{A}=\{\text{A,C,G,T}\}$. Standard phylogenetic models assume that each $x_{i}$ evolves independently according to a time-homogeneous continuous-time Markov chain (CTMC) defined by rate matrices $\mathbf{Q}^{(i)}$ with rates $\mathbf{Q}^{(i)}_{ab}=\gamma^{(i)}(b;a)$ where $\gamma^{(i)}:\mathscr{A}^{2}\rightarrow(0,+\infty)$ specifies the rate at which site $i$ mutates from $a$ to $b$. The independent-site assumption allows the straightforward calculation of sequence transition probabilities by
@@
3.2 Context-Dependent Mutation Models   The independent site assumption, while computationally convenient, fails to capture a variety of known effects which induce dependence between sites; examples include include CpG di-nucleotide mutability [Pederson:1998], structural constraints in RNA and proteins [Robinson:2003], and enzyme-driven somatic hypermutation in B-cell affinity maturation [Wiehe:2018, Mathews:2023b]. A variety of models have been introduced to account for these various effects, most of which can be written as context-dependent rate matrices [Mathews:2024]. For nucleotide sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{\ell})$, let $\tilde{x}_{i}=(x_{i-k},\ldots,x_{i},\ldots x_{i+k})$ denote the context of site $x_{i}$ for $i\in\{1,\ldots,\ell\}$. Relaxing the independent-site assumption, we can define a dependent site model (DSM) with sites evolving under a time-inhomogeneous CTMC and the rate at which $x_{i}$ mutates to $b\in\mathscr{A}$ in context $\tilde{x}_{i}$ given by
@@
3.3 Models of BCR Diversification for Lineage Reconstruction   As noted VDJ recombination plays a critical role in diversification, generating the initial gene segment rearrangement upon which SMH operates [hozumi_evidence_1976, tonegawa_somatic_1983]. Probabilistic hidden Markov models (HMMs) have been successfully applied to model VDJ recombination for clonal UCA reconstruction [kepler_reconstructing_2013, dhar_bayesian_2020]. These methods marry an independent-site mutation model (1) for SHM with an HMM for stochastic VDJ shuffling. In Section 4.6 below, we consider the incorporation of existing VDJ-models as prior distributions on the UCA, with a likelihood calculated from our dependent-site SHM model, using the methods described in the next section.
@@ 4.6 Incorporating VDJ Model-Based Priors on the UCA
Most applications of phylogenetic inference assume the common ancestor $\mathbf{x}$ is drawn from the stationary distribution of the evolutionary model. However, solving for the stationary distribution of a CTMC model with arbitrary site-dependence can be difficult, with the chain often non-reversible. (An exception is if the model satisfies conditions guaranteeing the stationary distribution is a Markov random field; see [jensen_probabilistic_2000].) For reconstructing BCR lineages, this assumption is not appropriate: the UCA is generated by VDJ recombination of germline sequences. Stochastic models of VDJ recombination are well-established [elhanati_inferring_2015, ralph_consistency_2016], and these models can provide informative prior distributions on the UCA during BCR lineage reconstruction [ralph_likelihood-based_2016, dhar_bayesian_2020]. Prior distributions on the UCA can be incorporated into the importance sampling approaches of Sections 4.3 and 4.4. Let $\zeta_{0}(\mathbf{x})$ denote the marginal distribution of $\mathbf{x}$ in the importance sampling proposal (e.g. the stationary distribution under ISM $p_{0}$), and $\zeta_{1}(\mathbf{x})$ the desired prior, then we have $p_{i}(Y,\mathbf{x},Z\mid g)=\zeta_{i}(\mathbf{x})\prod_{(\mathbf{u},\mathbf{v},t)\in E\left(g,\mathbf{x},Z,Y\right)}p_{i}(\mathbf{v}\mid\mathbf{u},t)$ and the importance weights can be calculated as before. However, for the VDJ model these priors may place mass on significantly different sequences, leading to inflation of importance weights. To prevent this, we instead we use the approach of [dhar_bayesian_2020] to sample from the posterior defined by the VDJ model combined with the ISM model for SHM: $\pi_{0,\text{VDJ}}\propto p_{0}(Y,Z\mid\mathbf{x},g)\zeta_{\text{VDJ}}(\mathbf{x})$. In this case the prior $\zeta_{\text{VDJ}}$ cancels out of the importance weights, and the estimator (13) remains unchanged.
Piecewise VDJ prior: In Section 5, we show that the VDJ prior improves reconstruction on germline-encoded regions, but sometimes causes errors in non-templated regions. We therefore define a piecewise prior under which the VDJ model applies outside the CDR3 region but not within the CDR3 itself. To draw from the posteriorwe use linearham [dhar_bayesian_2020] to sample $\pi_{0,\text{VDJ}}$ then substitute a CDR3 drawn from $p_{0}(\mathbf{x}_{\text{CDR3}}\mid Y,g)$ by pruning, resulting in the piecewise prior $p_{\text{VDJ}}(\mathbf{x}_{\overline{\mathrm{CDR3}}})\zeta_{0}(\mathbf{x}_{\mathrm{CDR3}})$, and the weight (10) remains correct.
@@ 5.2 Effects of context-dependence on tree and UCA reconstruction accuracy
Short Sequences: CDR3 regions The complementarity-determining region 3 (CDR3) of antibodies is a region of high variability that plays a crucial role in antigen recognition and binding. Therefore, we first focus on short sequences simulated from the CDR3 of a bnAb UCA. Two independent (tree, clonal sequence repertoire) sets were simulated. Samples from the posterior distribution over trees under the pure phylogenetic ISM model were obtained using Beast, and posterior probabilities under the DSM calculated from importance weights (13). Results for UCA reconstruction are summarized in Table 3a. The DSM consistently outperforms the ISM for all estimators. Also shown are UCAs reconstructed conditional on the true tree (unknown in practice); unsurprisingly, this significantly reduces UCA reconstruction error, but moreso for the ISM. Thus a key role of the DSM appears to be in improving the estimated tree.
@@
Full length BCR sequences Next we simulated 8 independent (tree, clonal sequence repertoire) sets of full length BCR sequences from an estimated UCA for HIV bnAb CH235. Each tree was pruned to select a subset of nodes to match individual Hamming distances of a real BCR dataset while ensuring no selected node was an ancestor/descendant of another, to approximate fixed-depth sequencing of BCR repertoires. Reconstruction accuracy is shown in Table 3a. In each simulated dataset, the true tree ranks as the highest probability under the DSM. For these longer sequences the number of reconstruction errors is larger, and the improvement obtained under the DSM is even clearer. (In Table 3 ISM refers to the mean-field DSM approximation (see Appdx 0.F); SHM to the pure phylogenetic model; VDJ to Linearham. IgPhyml does not do CDR3 junction inference [hoehn_repertoire-wide_2019] so does not provides a meaningful comparison.)

### v2

@@ Abstract
We introduce a method for approximating posterior probabilities of phylogenetic trees and reconstructing ancestral sequences under models of sequence evolution with site-dependence, where standard phylogenetic likelihood computations (pruning) fail. Our approach uses a combined data-augmentation and importance sampling scheme. A key advantage of our approach is the ability to leverage existing highly optimized phylogenetic software. We apply our approach to the reconstruction of B cell receptor affinity maturation lineages from high-throughput repertoire sequencing data and evaluate the impact of incorporating site-dependence on the reconstruction accuracy of both trees and ancestral sequences. We show that accounting for context-dependence during inference always improves the estimates of both ancestral sequences and lineage trees on simulated datasets. We also examine the impact of incorporating priors based on VDJ recombination models, and find that they significantly improve ancestral sequence reconstruction in germline-encoded regions, but increase errors in non-templated nucleotides. We propose a modified, piecewise prior to address this and demonstrate that it improves empirical reconstruction accuracy. We apply our approach to the analysis of the HIV broadly neutralizing antibodies DH270 and CH235 which are important targets of current vaccine design efforts.
@@ 1 Introduction
To prevent infectious diseases, the immune system must recognize and clear foreign pathogens capable of host infection. The adaptive immune system does so using B cells with receptors specifically tailored to bind pathogens. Facing a near-infinite variety of pathogen protein surfaces, B cells rely on two diversification methods: 1) initial gene segment shuffling (VDJ recombination) to enable weak binding to the diversity of antigens encountered, and 2) a refinement step (affinity maturation) that substantially improves binding and BCR-antigen recognition specificity, accomplished via somatic hypermutation (SHM). Understanding of this process and design of new vaccines is greatly aided by computational and statistical analysis of data from high-throughput B cell repertoire sequencing. A key challenge is accurate inference of BCR lineages and unmutated common ancestors, and significant progress has been made on these tools [kepler_reconstructing_2013, Ralph:2016, Marcou2018, dhar_bayesian_2020]. However, existing methods rely on conventional “independent-site" mutation models from molecular phylogenetics and unlike mutational processes elsewhere in the genome, these models do not adequately capture the SHM process, which is mediated by the enzyme activation induced cytidine deaminase (AID). Critically, AID activity is sequence-context dependent, targeting sites in certain favored motifs. Unfortunately, incorporation of context-dependent mutation models makes standard phylogenetic calculations intractable. Here we develop algorithms for reconstructing B cell lineages under context-dependent SHM models. We build on recent work on approximating pairwise sequence evolution probabilities [Mathews:2023b, Mathews:2025a, Mathews:2025b], extending these in non-trivial ways to trees using Monte Carlo methods of data augmentation, importance sampling, and sequential Monte Carlo. Notably, our approach can be implemented directly using existing phylogenetics software.
@@ 2.1 BnAbs, Sequential Vaccine Design, and B Cell Lineages
The goal of vaccination is to induce antibodies – secreted versions of B cell receptors (BCRs) – that neutralize pathogens and protect against infection. In infection, successive rounds of mutation and selection for antigen affinity (affinity maturation) enables B cells to recognize pathogens with specificity. However, rapidly mutating viruses (HIV, influenza, SARS-CoV-2) escape this B cell response by viral diversification. It is thus desired to elicit broadly neutralizing antibodies (bnAbs), but these are rarely elicited due to a variety of factors [Walker_Burton_2018]: low precursor frequencies in the human B cell receptor repertoire [Haynes_35962033, Jardine_27013733], highly mutated, and containing essential mutations that are highly improbable, creating acquisition bottlenecks [Wiehe:2018, Saunders_Wiehe_2019, Mathews:2023b]. Traditional vaccine design strategies have proven ineffective at eliciting bnAbs, motivating new strategies using known bnAbs as templates to design immunogens to guide B cell maturation the bnAb [Haynes_2012, Jardine_2015, Haynes_35962033, Steichen_2019, Kwong_29768174]. Sequential prime-boosting aims to first induce low-frequency bnAb precursors with a priming immunogen, then mature them with additional boosting immunogens, and relies heavily on inferring the precursor BCR sequence (the unmutated common ancestor (UCA)) from observed, clonally-related sequences [kepler_reconstructing_2013, Haynes_2012]. More generally reconstruction of B cell lineage - the set of affinity maturation pathways for clone members - provides insight into maturation bottlenecks and also precursors to serve as intermediate vaccine targets. Existing lineage reconstruction methods such as Cloanalyst [kepler_reconstructing_2013], IgOR [Marcou2018], Partis [Ralph:2016] and LinearHam [dhar_bayesian_2020] combine phylogenetic inference with models of VDJ recombination, but follow the vast majority of molecular phylogenetics in assuming that sequence positions evolve independently; this assumption is critical to tractability of phylogenetic inference. While appealing in their simplicity and computational convenience, these independent-site models fail to capture the critical context-dependence of SHM in affinity maturation.
@@ 2.2 Affinity Maturation of B Cell Receptors: Context-Dependent Somatic Hypermutation
S5F model and ARMADiLLO: We previously developed ARMADiLLO [Wiehe:2018] for simulating SHM using context-dependent mutation probabilities from the S5F model [Yaari2013], which captures context-dependence of SHM using context-dependent site-mutation and substitution probabilities derived from NGS data [Yaari2013] for the central nucleotide of 1,024 5-mer motifs, given its (2 upstream and 2 downstream) neighboring bases. The model parameterizes mutation probabilities as a product of mutability scores and substitution probabilities. ARMADiLLO has been used to forward-simulate SMH and as a null model to test for selection [Wiehe:2018, Tang:2023]. Unfortunately, this more realistic SHM model cannot be used in existing lineage reconstruction algorithms [kepler_reconstructing_2013, Marcou2018, Ralph:2016, dhar_bayesian_2020], as context-dependent mutation models directly violate the independent-site assumption that makes standard phylogenetic calculations (pruning) tractable. Since we might expect improved SHM models to provide improved reconstruction of UCAs and lineages, it is of great interest to develop methods for B cell lineage reconstruction which can handle context-dependent mutation models such as S5F/ARMADiLLO.
@@ 3 Background
3.1 Standard Models of DNA Evolution
Let $\mathbf{x}=(x_{1},\ldots,x_{\ell})$ and $\mathbf{y}=(y_{1},\ldots,y_{\ell})$ denote two nucleotide sequences of length $\ell$, where $x_{i},y_{i}\in\mathscr{A}$ for $\mathscr{A}=\{\text{A,C,G,T}\}$. Standard phylogenetic models assume that each $x_{i}$ evolves independently according to a time-homogeneous continuous-time Markov chain (CTMC) defined by rate matrices $\mathbf{Q}^{(i)}$ with rates $\mathbf{Q}^{(i)}_{ab}=\gamma^{(i)}(b;a)$ where $\gamma^{(i)}:\mathscr{A}^{2}\rightarrow(0,+\infty)$ specifies the rate at which site $i$ mutates from $a$ to $b$. The independent-site assumption allows the straightforward calculation of sequence transition probabilities by
@@ 3.1 Standard Models of DNA Evolution
3.2 Context-Dependent Mutation Models
The independent site assumption, while computationally convenient, fails to capture a variety of known effects which induce dependence between sites; examples include include CpG di-nucleotide mutability [Pederson:1998], structural constraints in RNA and proteins [Robinson:2003], and enzyme-driven somatic hypermutation in B-cell affinity maturation [Wiehe:2018, Mathews:2023b]. A variety of models have been introduced to account for these various effects, most of which can be written as context-dependent rate matrices [Mathews:2024]. For nucleotide sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{\ell})$, let $\tilde{x}_{i}=(x_{i-k},\ldots,x_{i},\ldots x_{i+k})$ denote the context of site $x_{i}$ for $i\in\{1,\ldots,\ell\}$. Relaxing the independent-site assumption, we can define a dependent site model (DSM) with sites evolving under a time-inhomogeneous CTMC and the rate at which $x_{i}$ mutates to $b\in\mathscr{A}$ in context $\tilde{x}_{i}$ given by
@@ 3.2 Context-Dependent Mutation Models
3.3 Models of BCR Diversification for Lineage Reconstruction
As noted VDJ recombination plays a critical role in diversification, generating the initial gene segment rearrangement upon which SMH operates [hozumi_evidence_1976, tonegawa_somatic_1983]. Probabilistic hidden Markov models (HMMs) have been successfully applied to model VDJ recombination for clonal UCA reconstruction [kepler_reconstructing_2013, dhar_bayesian_2020]. These methods marry an independent-site mutation model (1) for SHM with an HMM for stochastic VDJ shuffling. In Section 4.6 below, we consider the incorporation of existing VDJ-models as prior distributions on the UCA, with a likelihood calculated from our dependent-site SHM model, using the methods described in the next section.
@@ 4.6 Incorporating VDJ Model-Based Priors on the UCA
Most applications of phylogenetic inference assume the common ancestor $\mathbf{x}$ is drawn from the stationary distribution of the evolutionary model. However, solving for the stationary distribution of a CTMC model with arbitrary site-dependence can be difficult, with the chain often non-reversible. (An exception is if the model satisfies conditions guaranteeing the stationary distribution is a Markov random field; see [jensen_probabilistic_2000].) For reconstructing BCR lineages, this assumption is not appropriate: the UCA is generated by VDJ recombination of germline sequences. Stochastic models of VDJ recombination are well-established [elhanati_inferring_2015, ralph_consistency_2016], and these models can provide informative prior distributions on the UCA during BCR lineage reconstruction [Ralph:2016, dhar_bayesian_2020]. Prior distributions on the UCA can be incorporated into the importance sampling approaches of Sections 4.3 and 4.4. Let $\zeta_{0}(\mathbf{x})$ denote the marginal distribution of $\mathbf{x}$ in the importance sampling proposal (e.g. the stationary distribution under ISM $p_{0}$), and $\zeta_{1}(\mathbf{x})$ the desired prior, then we have $p_{i}(Y,\mathbf{x},Z\mid g)=\zeta_{i}(\mathbf{x})\prod_{(\mathbf{u},\mathbf{v},t)\in E\left(g,\mathbf{x},Z,Y\right)}p_{i}(\mathbf{v}\mid\mathbf{u},t)$ and the importance weights can be calculated as before. However, for the VDJ model these priors may place mass on significantly different sequences, leading to inflation of importance weights. To prevent this, we instead we use the approach of [dhar_bayesian_2020] to sample from the posterior defined by the VDJ model combined with the ISM model for SHM: $\pi_{0,\text{VDJ}}\propto p_{0}(Y,Z\mid\mathbf{x},g)\zeta_{\text{VDJ}}(\mathbf{x})$. In this case the prior $\zeta_{\text{VDJ}}$ cancels out of the importance weights, and the estimator (13) remains unchanged.
Piecewise VDJ prior:
In Section 5, we show that the VDJ prior improves reconstruction on germline-encoded regions, but sometimes causes errors in non-templated regions. We therefore define a piecewise prior under which the VDJ model applies outside the CDR3 region but not within the CDR3 itself. To draw from the posterior we use linearham [dhar_bayesian_2020] to sample $\pi_{0,\text{VDJ}}$ then substitute a CDR3 drawn from $p_{0}(\mathbf{x}_{\text{CDR3}}\mid Y,g)$ by pruning, resulting in the piecewise prior $p_{\text{VDJ}}(\mathbf{x}_{\overline{\mathrm{CDR3}}})\zeta_{0}(\mathbf{x}_{\mathrm{CDR3}})$, and the weight (10) remains correct.
@@ 5.2.1 Short Sequences: CDR3 regions
The complementarity-determining region 3 (CDR3) of antibodies is a region of high variability that plays a crucial role in antigen recognition and binding. Therefore, we first focus on short sequences simulated from the CDR3 of a bnAb UCA. Two independent (tree, clonal sequence repertoire) sets were simulated. Samples from the posterior distribution over trees under the pure phylogenetic ISM model were obtained using Beast, and posterior probabilities under the DSM calculated from importance weights (13). Results for UCA reconstruction are summarized in Table 3a. The DSM consistently outperforms the ISM for all estimators. Also shown are UCAs reconstructed conditional on the true tree (unknown in practice); unsurprisingly, this significantly reduces UCA reconstruction error, but moreso for the ISM. Thus a key role of the DSM appears to be in improving the estimated tree.
5.2.2 Full length BCR sequences
@@
Next we simulated 8 independent (tree, clonal sequence repertoire) sets of full length BCR sequences from an estimated UCA for HIV bnAb CH235. Each tree was pruned to select a subset of nodes to match individual Hamming distances of a real BCR dataset while ensuring no selected node was an ancestor/descendant of another, to approximate fixed-depth sequencing of BCR repertoires. Reconstruction accuracy is shown in Table 3a. In each simulated dataset, the true tree ranks as the highest probability under the DSM. For these longer sequences the number of reconstruction errors is larger, and the improvement obtained under the DSM is even clearer. (In Table 3 ISM refers to the mean-field DSM approximation (see Appdx 0.F); SHM to the pure phylogenetic model; VDJ to Linearham. IgPhyml does not do CDR3 junction inference [hoehn_repertoire-wide_2019] so does not provide a meaningful comparison.)
5.2.3 Improving UCA reconstruction with VDJ recombination models: